OpenStudy (jskinny0831):

In a recent stock market downturn, the value of a $5,000 stock is decreasing at 2.3% per month. This situation can be modeled by the equation A(t) = 5,000(0.977)12t, where A(t) is the final amount and t is time in years. Assuming the trend continues, what is the equivalent annual devaluation rate of this stock (rounded to the nearest tenth of a percent) and what is it worth (rounded to the nearest ten dollars) after 1 year? A. 24.4% and $3,780.00 B. 75.6% and $3,780.00 C. 27.6% and $1,380.00 D. 72.4% and $3,620.00

OpenStudy (jskinny0831):

do you know the answer? @AihberKhan

OpenStudy (aihberkhan):

Okay. This is the equation that is given to us: \(A(t) = 5,000(0.977){12t}\). All we need to do is substitute the \(1\), in for \(t\). This is because it says "after 1 year". So when we substitute the \(1\) in for \(t\), our equation should look like: \(A(t) = 5,000(0.977){12}\)

OpenStudy (jskinny0831):

yes and i got 3781

OpenStudy (jskinny0831):

but that would be A or B

OpenStudy (jskinny0831):

and when i have chose A it was wrong but now I'm just confused and scared to choose B

OpenStudy (aihberkhan):

It is more than \(2500\). This means that the annual rate of discount cannot be as much as 50%. Hmm... that is weird. It should be A. How about I tag someone else here to help you out, because I got A. Let me tag some people :)

OpenStudy (aihberkhan):

@Michele_Laino We are confused.... can you please help?

OpenStudy (aihberkhan):

@imqwerty We are confused.... can you please help?

OpenStudy (jskinny0831):

oh okay thank you because I'm so confused

OpenStudy (jskinny0831):

@mathmale

OpenStudy (michele_laino):

I think that the formula for \(A(t)\), is: \[\Large A\left( t \right) = 5000 \cdot {\left( {0.977} \right)^{12t}}\]

OpenStudy (jskinny0831):

yeah and i got 3781

OpenStudy (jskinny0831):

when you put 1 for the t

OpenStudy (aihberkhan):

Yeah thats what I thought @Michele_Laino .... but he wrote it differently so I got confused...

OpenStudy (aihberkhan):

Well the problem wrote it differently.

OpenStudy (michele_laino):

the requested percentage \(r\), is: \[\Large r = \frac{{A\left( 0 \right) - A\left( 1 \right)}}{{A\left( 0 \right)}} \cdot 100 = \frac{{5000 - 3782}}{{5000}} \cdot 100 = ...?\]

OpenStudy (jskinny0831):

24.4

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (jskinny0831):

omg i got it right!!

OpenStudy (jskinny0831):

i got 100%% but why did i get it wrong last time

OpenStudy (michele_laino):

I don't know

OpenStudy (michele_laino):

maybe the final result has been rounded in a wrong way

OpenStudy (jskinny0831):

can you guys help me with some other question i have to show my work

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

is the function like this one: \[\Large m\left( t \right) = 100 \cdot {0.5^{235t}}\]?

OpenStudy (jskinny0831):

its A(t)=100(0.5)^t/235

OpenStudy (jskinny0831):

if that makes sense

OpenStudy (michele_laino):

ok! then I rewrite it:\[\Large m\left( t \right) = 100 \cdot {0.5^{\left( {t/235} \right)}}\]

OpenStudy (jskinny0831):

t/235 is written like a fraction i guess thats the same but yeah

OpenStudy (michele_laino):

then for part A, we can note this: \[\Large {0.5^{\left( {t/235} \right)}} = {\left( {\frac{1}{2}} \right)^{\left( {t/235} \right)}} = \frac{1}{{{2^{\left( {t/235} \right)}}}} = {2^{ - \left( {t/235} \right)}}\]

OpenStudy (michele_laino):

so, after a substitution, we get: \[\huge m\left( t \right) = 100 \cdot {2^{ - \left( {t/235} \right)}}\]

OpenStudy (michele_laino):

for part B we have to compute this quantity: \[\Large m\left( {1000} \right) = 100 \cdot {2^{ - \left( {1000/235} \right)}} = ...{\text{grams}}\]

OpenStudy (jskinny0831):

is it 5.236 grams??

OpenStudy (michele_laino):

yes! Correct!!

OpenStudy (jskinny0831):

so i just put that for the answer in part B?

OpenStudy (jskinny0831):

and ook thank you!

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

:)

OpenStudy (jskinny0831):

okay i got it and can you help me with a couple more questions? :/

OpenStudy (jskinny0831):

The population of a particular city is given by the function P(t) = 12,500(1.04)^4t, where t is time in years and P(t) is the population after t years. Part A: Examine the function. What is the initial population and the percentage growth rate (rounded to the nearest whole percent)? Part B: What is the population size (rounded to the nearest whole person) in 10 years? Part C: Discuss collaboratively: How would the population size be affected if 0 < r < 1? Explain.

OpenStudy (michele_laino):

hint: the initial population is given by this quantity: \[\Large P\left( 0 \right) = 12500 \cdot {1.04^{\left( {4 \cdot 0} \right)}} = ...?\]

OpenStudy (jskinny0831):

yeah the t is the 0

OpenStudy (michele_laino):

yes!

OpenStudy (jskinny0831):

okay so what do i do for partA

OpenStudy (michele_laino):

what is \(P(0)=...?\)

OpenStudy (jskinny0831):

so i just write the function and put the 0 in for the T?

OpenStudy (michele_laino):

yes! The initial population is the value of \(P(t)\), when \(t=0\)

OpenStudy (jskinny0831):

ok :)thats what i put for part A

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} P\left( 0 \right) = 12500 \cdot {1.04^{\left( {4 \cdot 0} \right)}} = \hfill \\ \hfill \\ = 12500 \cdot {1.04^0} = 12500 \cdot 1 = ...? \hfill \\ \end{gathered} \]

OpenStudy (jskinny0831):

so 12500?

OpenStudy (michele_laino):

that's right!

OpenStudy (jskinny0831):

since multplying by 1 is the same

OpenStudy (jskinny0831):

so i put that ?

OpenStudy (michele_laino):

yes! It is the initial population

OpenStudy (michele_laino):

whereas, the percentage growth rate \(r\), I think it is given by the subsequent ratio: \[\huge r = \frac{{P\left( {t + 1} \right) - P\left( t \right)}}{{P\left( t \right)}} \cdot 100\]

OpenStudy (jskinny0831):

so whats that for? the percentage growth rate?

OpenStudy (michele_laino):

yes!

OpenStudy (jskinny0831):

oh okay yeh thats also part of part A

OpenStudy (jskinny0831):

so is it r=100/t

OpenStudy (michele_laino):

we have to substitute the function \(P(t)\): \[\huge P\left( t \right) = 12500 \cdot {1.04^{4t}}\] and after that substitution, we get: \[\Large \begin{gathered} r = \frac{{P\left( {t + 1} \right) - P\left( t \right)}}{{P\left( t \right)}} \cdot 100 = \hfill \\ \hfill \\ = \frac{{12500 \cdot {{1.04}^{4\left( {t + 1} \right)}} - 12500 \cdot {{1.04}^{4t}}}}{{12500 \cdot {{1.04}^{4t}}}} \cdot 100 = \hfill \\ \hfill \\ = \frac{{12500 \cdot {{1.04}^{4t}}\left( {{{1.04}^4} - 1} \right)}}{{12500 \cdot {{1.04}^{4t}}}} \cdot 100 = \hfill \\ \hfill \\ = \left( {{{1.04}^4} - 1} \right) \cdot 100 = ...\% \hfill \\ \end{gathered} \]

OpenStudy (jskinny0831):

16.98?

OpenStudy (jskinny0831):

%

OpenStudy (michele_laino):

yes! we have to round such value, so we get \(r=17 \%\)

OpenStudy (jskinny0831):

oh kook so ill put 17% thank you :')

OpenStudy (michele_laino):

:)

OpenStudy (jskinny0831):

got it so next is part B

OpenStudy (jskinny0831):

Part B-What is the population size (rounded to the nearest whole person) in 10 years?

OpenStudy (michele_laino):

here we have to compute this quantity: \[\huge P\left( {10} \right) = 12500 \cdot {1.04^{4 \times 10}} = ...?\]

OpenStudy (jskinny0831):

is it 6001.27?

OpenStudy (michele_laino):

I got: \(P(10)=60,012.75\)

OpenStudy (jskinny0831):

omg how?

OpenStudy (michele_laino):

we can write this: \[\Large \begin{gathered} P\left( {10} \right) = 12500 \cdot {1.04^{4 \times 10}} = \hfill \\ \hfill \\ = 12500 \cdot 4.80 = ...? \hfill \\ \end{gathered} \]

OpenStudy (jskinny0831):

oh jk i got it

OpenStudy (jskinny0831):

yeah i see what i did wrong

OpenStudy (michele_laino):

then we have to round such value: \[\Large \begin{gathered} P\left( {10} \right) = 12500 \cdot {1.04^{4 \times 10}} = \hfill \\ \hfill \\ = 12500 \cdot 4.80 \simeq 60013 \hfill \\ \end{gathered} \]

OpenStudy (jskinny0831):

yeah i just did that

OpenStudy (jskinny0831):

so thats the population size?

OpenStudy (michele_laino):

yes! It is the population after 10 years

OpenStudy (jskinny0831):

okok yay thank you !

OpenStudy (jskinny0831):

Part C: Discuss collaboratively: How would the population size be affected if 0 < r < 1? Explain.

OpenStudy (jskinny0831):

now i need this one

OpenStudy (michele_laino):

hint: if \[\Large 0 < r < 1\] then we can write: \[\Large \frac{{P\left( {t + 1} \right) - P\left( t \right)}}{{P\left( t \right)}} < 1\]

OpenStudy (michele_laino):

or more precisely: \[\Large 0 < \frac{{P\left( {t + 1} \right) - P\left( t \right)}}{{P\left( t \right)}} < 1\]

OpenStudy (jskinny0831):

so do i have to plug in some number for t right ?

OpenStudy (jskinny0831):

the r is supposed to go in the middle

OpenStudy (jskinny0831):

17%?

OpenStudy (michele_laino):

I think that we have to give a qualitative answer, namely we have to write a statement

OpenStudy (jskinny0831):

yeah we have to write how the population would be affected

OpenStudy (jskinny0831):

the population size

OpenStudy (michele_laino):

the condition \(0<r<1\) tells that the population is growing, since: \[\Large P\left( {t + 1} \right) - P\left( t \right) > 0 \Rightarrow P\left( {t + 1} \right) > P\left( t \right)\]

OpenStudy (jskinny0831):

so its affected because it is increasing in size

OpenStudy (michele_laino):

and on the other hand such growing is limited since we can write this: \[\Large \frac{{P\left( {t + 1} \right) - P\left( t \right)}}{{P\left( t \right)}} - 1 < 0 \Rightarrow P\left( {t + 1} \right) < 2P\left( t \right)\]

OpenStudy (michele_laino):

namely the population at a certain year doesn't exceed the double of the population of the preceding year

OpenStudy (jskinny0831):

so that would be the answer?

OpenStudy (jskinny0831):

that the population at a certain year doesn't exceed the double of the population of the preceding year?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (jskinny0831):

kook ill just write it and explain